Find ⋃ π‘¨π’Š ∞ π’Š=𝟏 and β‹‚ π‘¨π’Š ∞ π’Š=𝟏 where:


π‘¨π’Š = {π’Š,π’Š + 𝟏,π’Š + 𝟐, … } for every positive integer π’Š.


1
Expert's answer
2021-10-27T14:56:11-0400

Let us find "\cup_{i=1}^{\infty}A_i" and "\cap_{i=1}^{\infty}A_i" where "A_i = \{i,i+1,i+2, … \}" for every positive integer "i".


Taking into account that "A_i\supset A_j" for "j>i," we conclude that "\cup_{i=1}^{\infty}A_i=A_1=\{1,2,3,\ldots\}."


Let us show that "\cap_{i=1}^{\infty}A_i=\emptyset" using the method by contradiction. Suppose that "k\in \cap_{i=1}^{\infty}A_i" for some positive integer "k". Since "k\notin\{k+1,k+2,\ldots\}=A_{k+1}," we conclude that "k\notin\cap_{i=1}^{\infty}A_i." This contradiction proves that "\cap_{i=1}^{\infty}A_i=\emptyset."